OpenStudy (anonymous):

Find 12 of $45 Show and explain all your work 30 is what precent of 75?

OpenStudy (anonymous):

Is that 12% of $45??

OpenStudy (anonymous):

Yeah sorry

OpenStudy (anonymous):

So it will go like this: \[\frac{12}{100} \times 45\] can you solve this?

OpenStudy (anonymous):

5.4

OpenStudy (anonymous):

Yep..

OpenStudy (anonymous):

And my other ?

OpenStudy (anonymous):

Yes.. You are asked there what percentage?? So, let what percentage that we don't know is x: So this will become: \[\frac{x}{100} \times 75 = 30\] Can you here solve for x?

OpenStudy (anonymous):

Firstly multiply by 100 both the sides and tell me what you got?

OpenStudy (anonymous):

750 & 300

OpenStudy (anonymous):

How??

OpenStudy (anonymous):

100 x 75. 100 x 30

OpenStudy (anonymous):

100 X 75 = ?? and then: 30 X 100 = ??

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

So you got: \[\frac{x}{100} \times 75 \times 100 = 30 \times 100\] Here you can cancel 100 on Left Hand Side: \[\frac{x}{\cancel{100}} \times 75 \times \cancel{100} = 3000\] \[75x = 3000\] Now divide by 75 both the sides and find x..

OpenStudy (anonymous):

40

OpenStudy (anonymous):

Good..

OpenStudy (anonymous):

Thanks!! :)

OpenStudy (anonymous):

You are welcome dear..